trying to graft 3 points for -x-y=1

Answer by Alan3354(69443)

trying to graft 3 points for -x-y=1
-------
I think you mean graph.
---
-x - y = 1
x+y = -1
y = -x - 1
---
Pick values for x and find y.
Any 3 values